# Reason-RFT: Reinforcement Fine-Tuning for Visual Reasoning

### v1

@@ Abstract
Visual reasoning abilities play a crucial role in understanding complex multimodal data, advancing both domain-specific applications and artificial general intelligence (AGI). Existing methods improve VLM reasoning via Chain-of-Thought (CoT) supervised fine-tuning, using meticulously annotated training data to enhance visual reasoning capabilities. However, this training paradigm may lead to overfitting and cognitive rigidity, restricting the model’s ability to transfer visual reasoning skills across domains and limiting its real-world applicability. To address these limitations, we propose Reason-RFT, a novel reinforcement fine-tuning framework that significantly enhances generalization capabilities in visual reasoning tasks. Reason-RFT introduces a two-phase training framework for visual reasoning: (1) Supervised Fine-Tuning (SFT) with curated Chain-of-Thought (CoT) data activates the reasoning potential of Vision-Language Models (VLMs), followed by (2) Group Relative Policy Optimization (GRPO)-based reinforcement learning that generates multiple reasoning-response pairs, significantly enhancing generalization in visual reasoning tasks. To evaluate Reason-RFT’s visual reasoning capabilities, we reconstructed a comprehensive dataset spanning visual counting, structure perception, and spatial transformation, serving as a benchmark to systematically assess visual cognition, geometric understanding, and spatial generalization. Experimental results demonstrate Reasoning-RFT’s three key advantages: (1) Performance Enhancement: achieving state-of-the-art results across multiple tasks, outperforming most mainstream open-source and proprietary models; (2) Generalization Superiority: consistently maintaining robust performance across diverse tasks and domains, outperforming alternative training paradigms; (3) Data Efficiency: excelling in few-shot learning scenarios while surpassing full-dataset SFT baselines; Reason-RFT introduces a novel paradigm in visual reasoning, significantly advancing multimodal research. Project website: ReasonRFT.
Figure 1: Overview of Reason-RFT. Compared to traditional SFT-based methods, our proposed Reason-RFT framework demonstrates superior generalization in visual reasoning tasks, excelling in reasoning improvement, out-of-domain performance, and data efficiency.
footnotetext: Equal contribution.footnotetext: Project leaders.footnotetext: ${}^{\text{\Letter}}$ Corresponding author.
@@ 1 Introduction
Visual reasoning is pivotal for understanding complex multimodal data and advancing artificial general intelligence (AGI) [38, 45], making it a central focus in intelligent systems research. Recent advancements in image recognition [48, 27] and scene understanding [9, 64] have enabled transformative applications in healthcare [66, 47], robotics [28, 36], and autonomous driving [22, 20, 19, 21]. Consequently, enhancing visual reasoning capabilities has garnered significant attention from both industry and academia for its potential to drive transformative advancements.
@@
Researchers have explored two primary categories of methods to enhance visual reasoning capabilities: (1) neural-symbolic methods [13, 4, 7, 69, 16], which integrate symbolic reasoning with neural networks to improve interpretability and modularity, and (2) Supervised Fine-Tuning (SFT) based on vision-language models (VLMs) [63, 57], which utilize end-to-end training to strengthen reasoning abilities. However, both approaches face significant limitations. Neural-symbolic methods are hindered by high complexity and a strong reliance on program generation, while SFT is constrained by its dependence on high-quality Chain-of-Thought (CoT) annotated data and meticulously designed data mixing strategies, leading to issues such as overfitting, cognitive rigidity, and limited transferability from in-domain to out-of-domain tasks. These challenges reduce their effectiveness in real-world applications.
Recent advances in models such as GPT-o1 [45], DeepSeek-R1 [15], and Kimi-1.5 [56] have demonstrated that large-scale reinforcement learning (RL) during post-training significantly enhances reasoning abilities in mathematics and coding. By providing a dynamic and adaptive alternative to static datasets, RL effectively addresses traditional limitations and improves reasoning, even with minimal data. However, purely RL-based methods face challenges in generalizing across visual reasoning tasks, often limited by the model’s inherent cognitive constraints and the heterogeneity of multimodal data. Motivated by these observations, this paper investigates advanced RL techniques to enhance the adaptability and transferability of visual reasoning, aiming to overcome generalization challenges and improve performance in real-world applications.
@@
In this paper, we propose Reason-RFT, a novel two-phase reinforcement fine-tuning framework designed to enhance generalization in visual reasoning tasks. First, we employ Supervised Fine-Tuning (SFT) with Chain-of-Thought (CoT) reasoning to activate the model’s potential reasoning capabilities, using a high-quality domain-specific visual reasoning dataset tailored to stimulate related reasoning abilities. Subsequently, we further enhance reasoning potential through Group Relative Policy Optimization (GRPO), demonstrating that Reason-RFT achieves superior generalization by pushing the model’s reasoning limits. To evaluate its capabilities, we reconstructed a high-quality dataset spanning visual counting, structure perception, and spatial transformation, serving as a benchmark for assessing visual cognition, geometric understanding, and cross-task generalization. Extensive experiments highlight three key advantages of Reason-RFT: (1) Performance Improvement: It significantly outperforms mainstream VLMs in tasks like visual counting, structure perception, and spatial transformation; (2) Enhanced Generalization: It consistently surpasses SFT-only and RL-only paradigms across diverse tasks, validated through multi-dimensional evaluations; (3) Data Efficiency: It achieves over 95% of the performance of SFT-only approaches using less than 20% of the data. These results underscore the effectiveness and efficiency of Reason-RFT, offering a robust solution for advancing visual reasoning tasks.
Our main contributions are summarized as follows.
@@
  • We propose Reason-RFT, a novel two-phase reinforcement fine-tuning framework that significantly enhances the visual reasoning capabilities of VLMs by effectively combining the complementary strengths of SFT-based and RL-based methods.
  • We systematically analyze SFT-based and RL-based methods in visual reasoning tasks, and highlight the limitations of SFT-based approaches and the superior generalization of RL-based methods in reasoning improvement, out-of-domain performance, and data efficiency.
@@
  • We reconstructed a comprehensive dataset that covers three core domains: visual counting, structure perception, and spatial transformation, which serves as a benchmark to systematically assess visual cognition, geometric understanding, and cross-task generalization.
  • Extensive experiments validate the effectiveness of Reason-RFT, offering valuable insights to advance visual reasoning and introducing a novel paradigm that significantly pushes forward multi-modal research.
@@ 2 Related Work
Visual Reasoning Visual reasoning is a key challenge in advancing artificial general intelligence (AGI), requiring models to perform complex cognitive tasks based on visual perception [38, 42, 18]. It has broad applications, including visual counting [38, 37], geometric problem-solving [12, 30, 42, 70, 51], visual transformation reasoning [23], scientific research [41, 31], and robotic task planning [24, 28, 17]. Traditional methods rely on program generation [29, 16, 53] or neural-symbolic approaches [13, 4, 7, 69], while recent vision-language models (VLMs) leverage large language models (LLMs) for enhanced reasoning. For example, LLaVA-CoT [63] uses multi-stage supervised fine-tuning (SFT) with chain-of-thought (CoT) [61], and Insight-V [10] combines SFT with reinforcement learning (RL). DeepSeek-R1-Zero [14] introduces a rule-based RL method, significantly improving reasoning. Building on DeepSeek-R1 [14], our work compares SFT-based and RL-based paradigms, highlighting the superior performance of R1-based methods in visual reasoning.
Post-Training Post-Training is a crucial phase for enhancing the performance of LLMs and VLMs, bridging pre-trained models and their real-world applications [54, 68, 32, 8]. It primarily involves two methodologies: SFT [62, 58] and RL [71, 46, 52, 43, 65]. SFT adapts pre-trained models to specific tasks using task-oriented datasets, often formatted as instructions. Research like FLAN [60] highlights the importance of diverse instruction-tuning datasets for improving zero-shot performance, while iterative processes, such as Llama 3.1’s six-round strategy [11], integrate rejection sampling, synthetic data, and human annotations. RL aligns models with human preferences or task-specific goals through feedback mechanisms. Reinforcement Learning from Human Feedback (RLHF) [46] refines models using human preference data, as seen in Llama 3.1 [11] and Nemotron-4 [2], which use reward modeling techniques like DPO [49] and RPO [2]. For example, TÜLU 3 [34] employs length-normalized DPO, while DeepSeek-V3 [39] combines rule-based and model-based reward systems. Recently, DeepSeek-R1 [14] achieved significant text reasoning improvements through pure RL [50]. Our work adapts R1 methodologies to VLMs, enhancing visual reasoning, and systematically compares SFT-based and R1-based paradigms, demonstrating the superiority of R1-based methods in visual reasoning tasks.
Figure 2: Framework of Reason-RFT. Reason-RFT introduces a two-phase training framework for visual reasoning. First, Supervised Fine-Tuning (SFT) with Chain-of-Thought (CoT) reasoning activates the model’s reasoning potential using a high-quality domain-specific visual reasoning dataset in stage 1. Subsequently, in stage 2, Group Relative Policy Optimization (GRPO) enhances reasoning capabilities further, enabling Reason-RFT to achieve superior generalization by pushing the model’s reasoning limits. Specifically, reward evaluation consists of format reward and three different types of accuracy reward.
@@ 3 Methodology
3.1 Problem Definition
Visual reasoning can be formally defined as follows: given a visual input $I$ (e.g., images or videos) and a corresponding textual description or question $T$, the goal is to derive a conclusion or answer $A$ by analyzing the information in the visual input. This process can be represented as a mapping:
$R:(I,T)\rightarrow A,$
where $I\in\mathbb{R}^{H\times W\times C}$ is the visual input with $H$, $W$, and $C$ representing height, width, and channels, $T$ is the textual description or question typically in natural language, and $A$ is the derived conclusion or answer, often in natural language or structured data. Through this mapping, visual reasoning models integrate visual and textual information to achieve effective reasoning.
3.2 Reason-RFT
In this section, we propose Reason-RFT, a novel two-phase hybrid training strategy to enhance the reasoning capabilities of VLMs in complex visual reasoning tasks. As shown in Fig. 2, the framework comprises two stages: (1) SFT-based Visual Reasoning Activation, which uses SFT with high-quality CoT reasoning data to activate the model’s domain-specific reasoning capabilities, and (2) RL-based Visual Reasoning Reinforcement, which employs the GRPO algorithm with rule-based rewards to further push the upper limits of the model’s reasoning potential.
3.2.1 STAGE 1: SFT-based Reasoning Activation
In the initial phase, we employ Supervised Fine-Tuning (SFT) on a structured visual reasoning dataset containing step-by-step reasoning processes. This phase trains the model to decompose complex tasks into logical steps. Each sample is represented as $(x,q,r,a)$, where $x$ denotes the input images, $q$ is the question, $r$ is the reasoning steps, and $a$ is the final answer. The training objective maximizes the likelihood of generating both $r$ and $a$ given $(x,q)$:
@@ 3.2.1 STAGE 1: SFT-based Reasoning Activation
where $\mathcal{D}$ is the dataset, $y$ the concatenated sequence of $r$ and $a$, and $\pi_{\theta}$ the model’s token distribution. The output model $\pi_{\text{CoT}}$ serves as the initialization for the next stage, ensuring a robust foundation for reinforcement learning.
3.2.2 STAGE 2: RL-based Reasoning Enhancement
In the second phase, we refine $\pi_{\text{CoT}}$ using Group Relative Policy Optimization (GRPO), leveraging reinforcement learning for its efficiency and scalability. Unlike Proximal Policy Optimization (PPO), which requires a computationally expensive value network, GRPO calculates relative advantages by comparing rewards within a group of sampled actions, reducing computational overhead and simplifying optimization. This makes GRPO particularly suitable for visual reasoning tasks.
Sampling Action Groups For each input state $s=(x,q)$, where $x$ is the visual encoding of the input image and $q$ the textual encoding of the question, GRPO samples a group of actions $\{a_{1},a_{2},\dots,a_{G}\}$ from the current policy $\pi_{\theta}$, initialized from $\pi_{\text{CoT}}$. The sampling process is:
@@ 3.2.2 STAGE 2: RL-based Reasoning Enhancement
Reward Evaluation Each sampled action $a_{i}$ is assigned a reward $R(a_{i})$ based on verifiable criteria, resulting in a reward set $\{r_{1},r_{2},\dots,r_{G}\}$. For visual reasoning tasks, the reward function $R(a_{i})$ combines two components: Format Reward $R_{\text{format}}(a_{i})$ and Accuracy Reward $R_{\text{acc}}(a_{i})$. The Format Reward ensures structured response formats, while the Accuracy Reward evaluates correctness, balancing structured reasoning and factual accuracy. The reward function is defined as:
@@
Based on these advantages, the policy is updated to reinforce actions with positive advantages and reduce the probability of less effective ones. Policy updates are further constrained by minimizing the KL divergence between the updated and reference models, ensuring stable RL learning.
@@ 3.3 Reward Design for Visual Reasoning Tasks
For the diverse requirements of visual reasoning tasks, including Visual Counting, Structure Perception, and Spatial Transformation, our reward design integrates two essential components: Format Reward and Accuracy Reward. The Format Reward is uniformly applied across all tasks, ensuring that the model strictly adheres to a structured response format, which significantly enhances interpretability and consistency. For the Accuracy Reward, we carefully tailor the design to the specific characteristics of each task, as shown in Fig. 2, creating task-specific reward mechanisms to evaluate the correctness of the model’s responses.
Format Reward This component ensures structured and interpretable responses by requiring the model to adhere to a predefined template: reasoning within <think> and </think> and the final answer within <answer> and </answer>. A reward of 1 is given for strict adherence, while deviations result in a reward of 0.
@@
  • Discrete-valued Type This reward type applies to Visual Counting and partial Structure Perception tasks, where answers are discrete values (e.g., multiple-choice or integer-based responses). The accuracy reward $R_{\text{acc}}(a_{i})$ is defined as:
@@
    where $a_{\text{pred}}$ is the model’s predicted answer and $a_{\text{gt}}$ is the ground truth. This binary reward penalizes deviations from the exact ground truth, ensuring precision in tasks requiring unambiguous answers.
  • Mathematical Type This reward type is designed for Structure Perception tasks involving numerical answers, such as floating-point values or LaTeX-formatted expressions. It uses a tolerance-based evaluation to account for minor numerical deviations. The accuracy reward $R_{\text{acc}}(a_{i})$ is defined as:
@@
4 Experiment

### v3

@@ Abstract
Visual reasoning abilities play a crucial role in understanding complex multimodal data, advancing both domain-specific applications and artificial general intelligence (AGI). Existing methods improve Vision-Language Models (VLMs) reasoning via Chain-of-Thought (CoT) supervised fine-tuning, using meticulously annotated training data to enhance visual reasoning capabilities. However, this training paradigm may lead to overfitting and cognitive rigidity, restricting the model’s generalization ability to transfer visual reasoning skills under domain shift and limiting its real-world applicability. To address these limitations, we propose Reason-RFT, the first two-stage reinforcement fine-tuning framework for visual reasoning: (1) Supervised Fine-Tuning (SFT) with curated CoT data activates the reasoning potential of VLMs, followed by (2) Group Relative Policy Optimization (GRPO)-based reinforcement learning that generates multiple reasoning-response pairs, significantly enhancing the capability to address ubiquitous domain shift in visual reasoning tasks. To evaluate the visual reasoning capabilities of Reason-RFT, we reconstructed a comprehensive dataset encompassing visual counting, structural perception, and spatial transformation, serving as a benchmark for systematic assessment across three core dimensions. Experimental results demonstrate three key advantages: (1) Performance Enhancement: achieving state-of-the-art results across multiple tasks, outperforming mainstream open-source and proprietary models; (2) Generalization Superiority: consistently maintaining robust performance in addressing domain shift in typical visual reasoning tasks, outperforming alternative paradigms; (3) Data Efficiency: excelling in few-shot learning scenarios while surpassing full-dataset SFT baselines. Reason-RFT introduces a rebust training paradigm in visual reasoning, and please refer to project website: Reason-RFT.
footnotetext: Equal contribution. Project leader. $^{\text{\Letter}}$ Corresponding author: shanghang@pku.edu.cn.
@@ 1 Introduction
Visual reasoning is pivotal for understanding complex multimodal data and advancing artificial general intelligence (AGI) [1, 2], making it a central focus in intelligent systems research. Recent advancements in image recognition [3, 4, 5, 6, 7, 8, 9], interactive security [10, 11, 12] and scene understanding [13, 14] have enabled transformative applications in healthcare [15, 16], robotics [17, 18, 19, 20, 21, 22, 23, 24], and autonomous driving [25, 26, 27, 28, 29, 30, 31]. Consequently, enhancing visual reasoning capabilities has garnered significant attention from both industry and academia for its potential to drive transformative advancements.
@@
Researchers have explored two primary categories of methods to enhance visual reasoning capabilities: (1) neural-symbolic methods [32, 33, 34, 35, 36], which integrate symbolic reasoning with neural networks to improve interpretability and modularity, and (2) Supervised Fine-Tuning (SFT) based on vision-language models (VLMs) [37, 38], which utilize end-to-end training to strengthen reasoning abilities. However, both approaches face significant limitations. Neural-symbolic methods are hindered by high complexity and a strong reliance on program generation, while SFT is constrained by its dependence on high-quality Chain-of-Thought (CoT) annotated data and meticulously designed data mixing strategies, leading to issues such as overfitting, cognitive rigidity, and limited adaptability to domain shift. These challenges reduce their effectiveness in real-world applications.
Recent advances such as GPT-o1 [2], DeepSeek-R1 [39], and Kimi-1.5 [40] show that reinforcement learning (RL) during post-training enhances reasoning in coding and mathematics. RL offers a dynamic alternative to SFT by enabling exploration and feedback-driven optimization, which can improve performance with limited labeled data. However, pure RL methods often lack robustness to domain shifts—such as changes in visual appearance or configuration, limiting their generalization capacity in real-world visual reasoning scenarios.
Figure 1: Overview of Reason-RFT. Compared to traditional SFT-based methods, our proposed Reason-RFT framework demonstrates superior generalization in visual reasoning tasks, excelling in reasoning improvement, domain shift adaptability, and data efficiency.
@@
To address this, we propose Reason-RFT, the first two-stage reinforcement fine-tuning framework designed to enhance generalization in visual reasoning tasks. First, we employ SFT with CoT reasoning to activate the model’s potential reasoning capabilities, using a high-quality domain-specific visual reasoning dataset tailored to stimulate related reasoning abilities. Subsequently, we further enhance reasoning potential through Group Relative Policy Optimization (GRPO), demonstrating that Reason-RFT improves robustness under distribution shifts by enhancing the model’s reasoning capabilities. To evaluate its effectiveness, we constructed a high-quality dataset covering visual counting, structure perception, and spatial transformation, serving as a benchmark for evaluating three core capabilities of visual reasoning. Extensive experiments highlight three key advantages of Reason-RFT: (1) Performance Improvement: It achieves strong results on visual reasoning tasks, including visual counting, structure perception, and spatial transformation, outperforming mainstream VLMs; (2) Enhanced Generalization: It consistently exceeds both SFT-only and RL-only baselines under domain shift conditions, as demonstrated through comprehensive evaluations; (3) Data Efficiency: It reaches over 90% of the SFT-only performance while using less than 5% of the data. These results underscore the effectiveness and efficiency of Reason-RFT, establishing it as a robust framework for advancing visual reasoning. Our main contributions are summarized as follows.
@@
  • We introduce Reason-RFT, a two-stage reinforcement fine-tuning framework that significantly enhances the visual reasoning capabilities of VLMs by effectively combining the complementary strengths of SFT-based and RL-based methods.
  • We provide a systematic analysis of SFT-based and RL-based paradigms on visual reasoning tasks, identifying the limitations of SFT and the advantages of RL in improving reasoning ability, handling domain shifts, and achieving data-efficient learning.
@@
  • We reconstruct a comprehensive dataset spanning three core domains: visual counting, structure perception, and spatial transformation, serving as a benchmark for evaluating visual cognition, geometric understanding, and spatial generalization.
  • Extensive experiments validate the proposed framework, demonstrating its practicality and effectiveness, and providing a new perspective for reinforcement-driven multi-modal training.
@@ 2 Related Work
Visual Reasoning Visual reasoning is a core challenge in advancing AGI, requiring models to perform complex cognitive tasks grounded in visual perception [1, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. It underpins a wide range of applications, including visual counting [1, 51], geometric problem-solving [52, 53, 41, 54, 55], visual transformation reasoning [56, 57], scientific analysis [58, 59], and robotic task planning [60, 17, 61]. Traditional approaches rely on program generation [62, 36, 63] or neural-symbolic reasoning [32, 33, 34, 35], while recent advances in VLMs leverage large language models (LLMs) to enhance reasoning capabilities. For instance, LLaVA-CoT [37] employs multi-stage SFT with CoT prompting [64], and Insight-V [65] integrates SFT with RL. DeepSeek-R1-Zero [66] further introduces a rule-guided RL framework that substantially improves reasoning performance. Building upon the DeepSeek-R1 [66], our work provides a comparative analysis of SFT-based and RL-based paradigms, demonstrating the advantages of R1-style methods in enhancing visual reasoning.
Post-Training Post-Training is a crucial phase for enhancing the performance of LLMs and VLMs, bridging pre-trained models and their real-world applications [67, 68, 69, 70]. It primarily involves two methodologies: SFT [71, 72] and RL [73, 74, 75, 76, 77]. SFT adapts pre-trained models to specific tasks using task-oriented datasets, often formatted as instructions. Research like FLAN [78] highlights the importance of diverse instruction-tuning datasets for improving zero-shot performance, while iterative processes, such as Llama 3.1’s six-round strategy [79], integrate rejection sampling, synthetic data, and human annotations. RL aligns models with human preferences or task-specific goals through feedback mechanisms. Reinforcement Learning from Human Feedback (RLHF) [74] refines models using human preference data, as seen in Llama 3.1 [79] and Nemotron-4 [80], which use reward modeling techniques like DPO [81] and RPO [80]. For example, TÜLU3 [82] employs length-normalized DPO, while DeepSeek-V3 [83] combines rule-based and model-based reward systems. Recently, DeepSeek-R1 [66] achieved significant text reasoning improvements through pure RL [84]. Our work first adapts R1 methodologies to VLMs, enhancing visual reasoning, and systematically compares SFT-based and RL-based paradigms in visual reasoning tasks.
@@ 3 Methodology
In this section, we introduce Reason-RFT, a novel two-stage training strategy to enhance the reasoning capabilities of VLMs in complex visual reasoning tasks. As shown in Fig. 2, the framework comprises two stages: (1) SFT-based Visual Reasoning Activation, which uses SFT with high-quality CoT reasoning data to activate the model’s domain-specific reasoning capabilities, and (2) RL-based Reasoning Enhancement, which employs the GRPO algorithm with rule-based rewards to further push the upper limits of the model’s reasoning potential.
3.1 STAGE 1: SFT-based Reasoning Activation
In the initial stage, we employ SFT on a structured visual reasoning dataset containing step-by-step reasoning processes. This stage trains the model to decompose complex tasks into logical steps. Each sample is represented as $(x,q,r,a)$, where $x$ denotes the input images, $q$ is the question, $r$ is the reasoning steps, and $a$ is the final answer. The training objective maximizes the likelihood of generating both $r$ and $a$ given $(x,q)$:
@@ 3.1 STAGE 1: SFT-based Reasoning Activation
where $\mathcal{D}$ denotes the dataset, $y$ represents the concatenated sequence of $r$ and $a$, and $\pi_{\theta}$ denotes the model’s token distribution. The resulting model $\pi_{\text{CoT}}$ is used to initialize the subsequent stage, providing a stable foundation for RL-based reasoning enhancement.
3.2 STAGE 2: RL-based Reasoning Enhancement
In the second stage, we refine $\pi_{\text{CoT}}$ using GRPO, leveraging RL for its efficiency and scalability. Unlike Proximal Policy Optimization (PPO), which requires a computationally expensive value network, GRPO calculates relative advantages by comparing rewards within a group of sampled actions, reducing computational overhead and simplifying optimization. This makes GRPO particularly suitable for visual reasoning tasks.
Sampling Action Groups For each input state $s=(x,q)$, GRPO samples a group of actions $\{a_{1},a_{2},\dots,a_{G}\}$ from the current policy $\pi_{\theta}$, initialized from $\pi_{\text{CoT}}$. The sampling process is:
@@ 3.2 STAGE 2: RL-based Reasoning Enhancement
Reward Evaluation. Each sampled action $a_{i}$ receives a reward $R(a_{i})$ based on verifiable criteria, yielding a reward set $\{r_{1},r_{2},\dots,r_{G}\}$. For visual reasoning tasks, the reward $R(a_{i})$ is composed of a format reward $R_{\text{format}}(a_{i})$, which enforces structured outputs, and an accuracy reward $R_{\text{acc}}(a_{i})$, which measures correctness. This formulation balances structural alignment and factual precision in reasoning. The reward function is defined as:
@@
Based on these advantages, the policy is updated to reinforce actions with positive advantages and reduce the probability of less effective ones. To maintain training stability, the update is constrained by minimizing the KL divergence between the updated and reference policies.
Figure 2: Framework of Reason-RFT. Reason-RFT adopts a two-stage training paradigm for visual reasoning. The first stage applies SFT with CoT reasoning to establish strong task-specific priors. In the second stage, GRPO is employed to further enhance reasoning capability and generalization.
@@ 3.3 Reward Design for Visual Reasoning Tasks
For the diverse requirements of visual reasoning tasks, including visual counting, structure perception, and spatial transformation, our reward design integrates two essential components: format reward and accuracy reward. The format reward is uniformly applied across all tasks, ensuring that the model strictly adheres to a structured response format for consistency. For the accuracy reward, we carefully tailor the design to the specific characteristics of each task, as shown in Fig. 2, creating task-specific reward mechanisms to evaluate the correctness of the model’s responses.
Format Reward This component ensures structured and interpretable responses by requiring the model to adhere to a predefined template: reasoning within <think> and </think> and the final answer within <answer> and </answer>. A reward of 1 is only given for strict adherence.
@@
  • Discrete-valued Type This reward type applies to visual counting and structure perception tasks, where answers are discrete values (e.g., multiple-choice or integer-based responses). The accuracy reward $R_{\text{acc}}(a_{i})$ is defined as:
@@
    where $a_{\text{pred}}$ is the predicted answer and $a_{\text{gt}}$ is the ground truth. This discrete reward penalizes deviations from ground truth, ensuring precision in tasks requiring unambiguous answers.
  • Mathematical Type This reward type is designed for structure perception tasks involving numerical answers, such as floating-point values or LaTeX-formatted expressions. It uses a tolerance-based evaluation to account for minor numerical deviations, which is defined as:
@@
    where $T_{\text{pred}}^{f+o+v}$ is the subset of transformation steps with complete matches (w/ function, object, and value), $T_{\text{pred}}^{f+o/v}$ are the subsets with partial and only-function matches (w/ function and object, or w/ function and value), $T_{\text{pred}}^{f}$ is the subset with only-function matches. $\alpha$ and $\beta$ are the weighting coefficients for partial matches. This formulation ensures nuanced evaluation for flexible adjustment of partial match contributions.
4 Experiments
We design experiments to investigate the following key research questions:
  • RQ1: How effective is Reason-RFT in reasoning, generalization, and data efficiency?
  • RQ2: Why is the STAGE 1 of SFT with CoT reasoning necessary?
  • RQ3: Why is the STAGE 2 of reinforcement fine-tuning necessary?
  • RQ4: How does reward design affect Reason-RFT’s performance?
  • RQ5: What training dynamics emerge during reinforcement fine-tuning, and how do they shape the reasoning behavior of Reason-RFT?